OpenStudy (anonymous):

Simplify. A. 11 B. 13 C. 14 D. 25

OpenStudy (anonymous):

Simplify \[3^{2} + 2^{4}\]

OpenStudy (dtan5457):

3^2=? 2^4=? Add the two values together

OpenStudy (anonymous):

is that 3X2 and 2X4?

OpenStudy (anonymous):

@Rocky_ 3 x 2 = 6 2 x 4 = 8 6 + 8 = ?

OpenStudy (dtan5457):

@PizzaLover123 3^2=3 x 3 2^4= 2 x 2 x 2 x 2

OpenStudy (anonymous):

well, do you know what powers are ? it is multiplying the whole number by itself as many times as the number obove says. for example, 5^2 = 25 5 * 5 = 25, and that how you answer. 3^2 3 * 3 = 9 so, what is 2^4 ? could you do that ?

OpenStudy (anonymous):

Wait no let me see

OpenStudy (anonymous):

Never mind its 25 Thanks @dtan5457

OpenStudy (anonymous):

don't give out answers, @PizzaLover123

OpenStudy (anonymous):

Ok thanks i think i got it?

OpenStudy (anonymous):

@silverr thanks for the help..

OpenStudy (anonymous):

okay, @Rocky_ ! and okay, @PizzaLover123

OpenStudy (anonymous):

@silverr Im kindof new here :p But yea I gave you a badge :) or medal

OpenStudy (anonymous):

that's okay ! i did it before too ( i did it recently, actually, because i forgot )